# Extracting Information in a Low-resource Setting: Case Study on Bioinformatics Workflows

### v1

@@ 1 Introduction
To overcome the low-resource challenges in the field, several strategies can be employed. First, decoder-based approaches, such as those using generative models, offer a potential solution, although they may not always outperform other methods in terms of extraction accuracy [3, 17]. Next, encoder-based approaches can be explored in two ways: using larger, related corpora to enrich model training, or by developing a small, domain-specific training corpus to focus on the extraction of entities relevant to the field. Finally, another promising approach is the injection of domain-specific knowledge directly into language models.
The challenge is thus to develop a simple and automated approach capable of extracting information about workflows described in the literature, aiming to not only systematically document workflows presented in the literature but also facilitate their registration in workflow repositories. This publication introduces a methodology for extracting information about workflows from the literature.
The contributions of this work are as follows: (a) a representation of workflow components using a schema comprising 16 entities and a corpus manually annotated named BioToFlow; (b) experiments with few shot Named-Entity Recognition (NER) using an autoregressive language model; (c) NER experiments using a pre-existing corpus and BioToFlow and (d) integration of workflow knowledge into NER models.
@@ 2.1 Workflow Representation Schema
To systematically describe the bioinformatics workflows and their components, we relied on discussions with workflow experts (notably from the ELIXIR and BOSC International Bioinformatics communities) and the review of several articles describing workflows (e.g., [14, 25, 4, 5, 15]). Generally, bioinformatics workflows consist of data analysis steps, each containing a script that may or may not invoke a bioinformatics tool. For a workflow to be executed, it is crucial to maintain a record of its execution environment. Figure 1 presents the proposed schema, which distinguishes between three main categories of entities. The first category is core entities, representing major components of a workflow, such as bioinformatics tools and data. Tools in this category can also be further classified based on their context: whether they are explicitly identified as bioinformatics tools (BioInfo), laboratory-related tools (Lab), tools mentioned but not used in the final workflow (Context), or tools referenced in a general sense without specific details (General). The second category includes environment entities, which capture the resources and configurations required for workflow execution, such as the execution environment or programming languages used in scripts. Finally, the third category addresses more specific details, including versioning information, bibliographic references describing tools or libraries, as well as descriptions and parameters associated with certain elements.
@@ 2.2.1 Article Selection Criteria
Articles describing bioinformatics workflows with a link to workflow code in Github have been targeted using a query on PubMed and PubMedCentral [19] [^1] PubMed query: (nextflow[Abstract] OR snakemake[Abstract] OR nextflow[Title] OR snakemake[Title]) AND github[All Fields]. As of 11/18/2024, it returns 124 and 123 articles describing workflows from Nextflow and Snakemake systems respectively.
@@ 2.2.2 Creation of an Annotated Corpus
To validate this information representation schema and create a resource for evaluating the extraction of such information, we developed an annotated corpus using a collaborative method [7] with the help of an annotation guide that we designed. Annotations were performed using BRAT (Brat Rapid Annotation Tool [21], v1.3 p1), recognized for its qualities including the ability to process nested entities, an existing annotation schema, and pre-annotations [18].
@@
Overall, our corpus, named BioToFlow [^3] https://doi.org/10.5281/zenodo.11204427, is a small corpus composed of 52 articles (26 articles related to Nextflow workflows and 26 on Snakemake workflows) with a total of 78,419 tokens (27,786 annotated tokens). However, the entities are diversified and distributed according to Table 1. In our corpus, we allow the presence of nested entities ($\sim$8of entities in BioToFlow; an example can be seen in Figure 2 where a WorkflowName entity is nested within a Biblio entity).
@@ 3 Decoder-based Approach for Named Entity Recognition
Because of the small size of the BioToFlow corpus, we sought to leverage an auto-regressive language model to perform few-shot named entity recognition. Specifically, we used Llama-3-8B-Instruct, which demonstrated the best performance among generative models in a comparative study of masked language models and generative models utilizing prompting techniques [17]. Following the same methodology as in this study, we conducted a series of few-shot experiments, varying the number of examples provided and testing different prompt formulations. We use random sampling to split the BioToFlow corpus: we use 75% of the corpus (39 articles) for training and the remaining 25% (13 articles) for testing. The results obtained on the BioToFlow test set using the Llama-3-8B-Instruct model with the best hyperparameters are presented in Table 2.
@@
We obtained an overall performance below 40% in relaxed mode and 31.2% in strict mode (not shown). These results are consistent with findings on other biomedical corpora [17] and lead us to explore alternative methods leveraging encoder-based models.
@@ 4.1 Experiments on a Large Corpus: SoftCite
For the extraction of our target entities, we chose a biLSTM-CRF neural model [24], implemented by the NLStruct Python library package [^4] https://github.com/percevalw/NLStruct (v0.2.0). This library offers the dual advantage of supporting the detection of nested entities, which is essential for certain entity types in our dataset, and accepting input files in the BRAT format. Additionally, it has demonstrated strong performance in the biomedical domain, which closely aligns with our field. More specifically, we used this model with the SciBERT language model [2], more specifically trained on scientific articles. We conventionally chose to randomly divide the articles into two main sets: the first set, 75% of the corpus (927 articles), was used to create the training set, while the remaining 25% (232 articles) constituted the test set. Within the training set, two-thirds of the articles (649 articles) were used for the actual training of the models and one-third (278 articles) for their validation. A distinctive aspect of our approach is that we constructed five splits through random sampling to minimize dependency on a specific division. For each split, five versions of the entity recognition model were produced using different random seeds, ensuring more robust and reproducible results.
@@ 4.2 Experiments on a Small Corpus: BioToFlow
The first results show that training a model on our little corpus obtains higher results for all the entities (see Table 2, 5, and 6).
@@ 5.0.2 Adding Vocabulary
To improve the extraction of bioinformatics tool names, another strategy involves incorporating external knowledge into the encoder. Since the knowledge to inject in our case is made of lists of names, we chose to perform this injection at the level of its vocabulary, similarly to work such as [11]. Various knowledge bases of bioinformatics tools exist, from which tool and binary names can be extracted. We utilized four knowledge bases: Biotools [12], Bioconda [10], Biocontainers [22] and Bioweb [^5] https://bioweb.pasteur.fr/welcome. From these knowledge bases, we aggregated the names of tools and binaries to create an enriched vocabulary. This new vocabulary was then injected into the SciBERT language model to improve its ability to recognize bioinformatics tool names. We evaluated this strategy by conducting two experiments on BioToFlow, testing two approaches: (a) fine-tuning the SciBERT model with the new vocabulary and (b) using the enriched vocabulary without additional fine-tuning. In both cases, the embeddings of the added tokens were initialized by averaging the embeddings of their subtokens.
@@ 6 Conclusion
Based on a workflow representation schema, we developed a small entity-rich corpus, BioToFlow. Using this corpus, we experimented with various entity extraction techniques, including decoder-based models, which achieved scores below 50%, and encoder-based methods. These included leveraging a large corpus with annotations covering a subset of the schema and our BioToFlow corpus. A final approach sought to incorporate additional knowledge into the models. The most effective strategy involved training a model on our small corpus while integrating a vocabulary list specific to bioinformatics tools, which improved the extraction of this entity type. To extend this latest approach, we will investigate various ways to enhance the embeddings of the added vocabulary, starting with their adaptation during the training of the NER models, similarly to [11]. More globally, the entity linking between the components extracted from workflows described in scientific literature and those found in code from public repositories is an essential task that we will consider to merge information from both sources.
@@ 6.0.1 Acknowledgements
We sincerely thank N. Bossut, A. Gaignard, G. Marchment, M. Schmit, H. Menager, F. Lemoine for their help with the annotation process. This work has received support from the French government (Agence Nationale pour la Recherche) under the France 2030 program grant agreement ANR-22-PESN-0007 (ShareFAIR).

### v2

@@ 1 Introduction
To overcome the low-resource challenges in the field, we consider in this paper several strategies. First, decoder-based approaches, such as those using generative models, offer a potential solution, although they may not always outperform other methods in terms of extraction accuracy [3, 18]. Next, encoder-based approaches can be explored in two ways: using larger, related corpora to enrich model training, or by developing a small, domain-specific training corpus to focus on the extraction of entities relevant to the field. Finally, another promising approach is the injection of domain-specific knowledge directly into language models.
From a methodological perspective, we address in this paper several research questions related to low-resource settings. In a specific domain, what resources should be annotated, in what quantity and how should domain experts be involved? Which existing resources can be reused or adapted for the new task? What methods are the most effective for the target task and how can they be combined? From a more application-oriented perspective, the challenge is to develop a simple and automated approach capable of extracting information about workflows described in the literature, aiming to not only systematically document workflows presented in the literature but also facilitate their registration in workflow repositories.
In preliminary work [22], we presented a corpus of scientific articles annotated with entities representing bioinformatics workflow information (BioToFlow). We also reported preliminary experiments with Named-Entity Recognition (NER) from the corpus with a specific focus on the memorization vs. generalization abilities of statistical and rule-based methods. Herein, we provide a brief description of BioToFlow and we present new contributions as follows: (a) experiments with few-shot NER using an autoregressive language model; (b) NER experiments using a pre-existing corpus and BioToFlow and (c) integration of workflow knowledge into NER models. Our code is publicly available [^1] https://gitlab.liris.cnrs.fr/sharefair/ner_workflows.
@@ 2.1 Workflow Representation Schema
To systematically describe the bioinformatics workflows and their components, we relied on discussions with workflow experts (notably from the ELIXIR and BOSC International Bioinformatics communities) and the review of several articles describing workflows (e.g., [4, 5, 15, 16, 27]. Generally, bioinformatics workflows consist of data analysis steps, each containing a script that may or may not invoke a bioinformatics tool. For a workflow to be executed, it is crucial to maintain a record of its execution environment. Figure 1 presents the proposed schema, which distinguishes between three main categories of entities. The first category is core entities, representing major components of a workflow, such as the name of the workflow described, bioinformatics tools, and data. Tools can also be further classified based on the context: whether they are explicitly identified as bioinformatics tools (BioInfo), laboratory-related tools (Lab), tools mentioned but not used in the final workflow (Context), or tools referenced in a general sense without specific details (General). One entity is dedicated to methods, i.e. the descriptions of workflow steps, but without mentioning specific tool names (e.g., "mapping DNA sequences against a genome" vs. "BWA"). The data refers to the information processed by the workflow (e.g., "DNA sequence"). Articles exhibit different levels of specificity and can contain mentions of a database name or the exact filenames that were used. However, in other cases, they remain more general and only refer to broad categories of data. The second category includes environment entities, which capture the resources and configurations required for workflow execution, such as the execution environment (e.g., the management system used, the hardware it runs on) or programming languages used in scripts. Finally, the third category addresses more specific details, including versioning information, bibliographic references describing tools or libraries, as well as descriptions and parameters associated. This additional information can be associated with the entities present in the workflow body or be more closely related to the workflow environment, providing further context about it.
@@ 2.2.1 Article Selection Criteria
Articles describing bioinformatics workflows with a link to workflow code in Github have been targeted using a query on PubMed and PubMedCentral [20] [^2] PubMed query: (nextflow[Abstract] OR snakemake[Abstract] OR nextflow[Title] OR snakemake[Title]) AND github[All Fields]. During the annotation phase (February 2024), this query returned 89 and 91 articles describing workflows from Nextflow and Snakemake systems respectively, from which we randomly selected articles for annotation. As of 11/18/2024, the same query now returns 124 and 123 articles, demonstrating a continuous increase in the number of published articles on these workflows. Articles were pre-processed to select the contents of the Material and Methods and Implementation sections, which commonly contain workflow descriptions that we aim to annotate.
@@ 2.2.2 Creation of an Annotated Corpus
To validate the information representation schema described in section 2.1 and create a resource for evaluating the extraction of such information, we developed an annotated corpus using a collaborative method [7] with the help of an annotation guide that we designed. Annotations were performed using BRAT (Brat Rapid Annotation Tool [23], v1.3 p1), recognized for its qualities including the ability to process nested entities, an existing annotation schema, and pre-annotations [19].
@@
Overall, our corpus, named BioToFlow [^4] https://doi.org/10.5281/zenodo.14900544, is a small corpus composed of 52 articles (26 articles related to Nextflow workflows and 26 on Snakemake workflows, randomly selected from PubMed while ensuring a balanced representation of both workflow management systems) with a total of 78,419 tokens (27,786 annotated tokens). However, the entities are diversified and distributed according to Table 1. In our corpus, we allow the presence of nested entities ($\sim$8of entities in BioToFlow; an example can be seen in Figure 2, where a WorkflowName entity is nested within a Biblio entity).
@@ 3 Decoder-based Approach for Named Entity Recognition
Because of the small size of the BioToFlow corpus, we sought to leverage an auto-regressive language model to perform few-shot named entity recognition. Specifically, we used Llama-3-8B-Instruct, which demonstrated the best performance among generative models in a comparative study of masked language models and generative models utilizing prompting techniques [18]. Following the same methodology as in this study [^5] The code is available at https://github.com/marconaguib/autoregressive_ner, we conducted a series of few-shot experiments, varying the number of examples provided (either 5 or 10) and testing different prompt formulations, such as providing or omitting entity definitions and task framing.
We use random sampling to split the BioToFlow corpus: we use 75% of the corpus (39 articles) for training and the remaining 25% (13 articles) for testing. The results obtained on the BioToFlow test set using the Llama-3-8B-Instruct model with the best hyperparameters are presented in Table 2.
@@
We obtained an overall performance below 40% in relaxed mode and 31.2% in strict mode (not shown). These results are consistent with findings on other biomedical corpora [18]. Bioinformatics workflows are complex, and although they are closely related to the medical domain, the entities are often very specific. This leads us to explore alternative methods leveraging encoder-based models.
@@ 4.1 Experiments on a Large Corpus: SoftCite
For the extraction of our target entities, we chose a biLSTM-CRF neural model (combination of a bidirectional LSTM (for sequence processing) with CRF (for prediction)) [26], implemented by the NLStruct Python library package [^6] https://github.com/percevalw/NLStruct (v0.2.0), which performs multiclass classification. All the experiments with NLStruct were run with the hyperparameters provided in Appendix 0.A. This library offers the dual advantage of supporting the detection of nested entities, which is essential for certain entity types in our dataset, and accepting input files in the BRAT format. Additionally, it has demonstrated strong performance in the biomedical domain, which closely aligns with our field. More specifically, we used this model with the SciBERT language model [2], more specifically trained on scientific articles. We conventionally chose to randomly divide the articles into two main sets: the first set, 75% of the corpus (927 articles), was used to create the training set, while the remaining 25% (232 articles) constituted the test set. Within the training set, two-thirds of the articles (649 articles) were used for the actual training of the models, and one-third (278 articles) for their validation. A distinctive aspect of our approach is that we constructed five splits through random sampling to minimize dependency on a specific division. For each split, five versions of the entity recognition model were produced using different random seeds, ensuring more robust and reproducible results.
@@ 4.2 Experiments on a Small Corpus: BioToFlow
The first results show that training a model on our little corpus obtains higher results for all the entities (see Tables 2, 5, and 6).
@@ 5.0.2 Adding Vocabulary
Tools are considered a key entity in our work because each tool has a specific function in the workflow, and it is essential for the overall understanding of the workflow task that all tools are extracted properly. To improve the extraction of bioinformatics tool names, another strategy involves incorporating external knowledge into the encoder. Since the knowledge to inject in our case is made of lists of names, we chose to perform this injection at the level of its vocabulary, similarly to work such as [9, 12]. Various knowledge bases of bioinformatics tools exist, from which tool and binary names (the executable files or commands used to run the tools in a program) can be extracted. We utilized four knowledge bases: Biotools [13], Bioconda [11], Biocontainers [24] and Bioweb [^7] https://bioweb.pasteur.fr/welcome. From these knowledge bases, we aggregated the names of tools and binaries to create an enriched vocabulary. This new vocabulary was then injected into the SciBERT language model to improve its ability to recognize bioinformatics tool names. We evaluated this strategy by conducting two experiments on BioToFlow, testing two approaches: (a) fine-tuning the SciBERT model with the new vocabulary and (b) using the enriched vocabulary without additional fine-tuning. In both cases, the embeddings of the added tokens were initialized by averaging the embeddings of their subtokens.
@@ 6 Conclusion
Based on a workflow representation schema, we developed a small entity-rich corpus, BioToFlow. Using this corpus, we experimented with various entity extraction techniques, including decoder-based models, which achieved scores below 50%, and encoder-based methods. These included leveraging a large corpus with annotations covering a subset of the schema and our BioToFlow corpus. A final approach sought to incorporate additional knowledge into the models. The most effective strategy involved training a model on our small corpus while integrating a vocabulary list specific to bioinformatics tools, which improved the extraction of this entity type. All these experiments and combinations of approaches demonstrate that bioinformatics workflows represent a highly specialized branch of biology. This implies that named entity extraction requires a large amount of domain-specific annotated data, as well as the use of language models tailored to bioinformatics. To improve entity extraction, additional few-shot learning approaches should be explored, such as using NER models trained to identify any type of entity, like GLiNER [28]. To further extend the approach of incorporating domain knowledge, we will also investigate various ways to enhance the embeddings of the added vocabulary, starting with their adaptation during the training of the NER models, similarly to [12]. More globally, the entity linking between the components extracted from workflows described in scientific literature and those found in code from public repositories is an essential task that we will consider to merge information from both sources.
@@ 6.0.1 Acknowledgements
We sincerely thank N. Bossut, A. Gaignard, G. Marchment, M. Schmit, H. Menager, F. Lemoine for their help with the annotation process. We also thank M. Naguib for his help with the decoder-based experiments. This work has received support from the French government (Agence Nationale pour la Recherche) under the France 2030 program grant agreement ANR-22-PESN-0007 (ShareFAIR).
@@
Appendix 0.A Hyperparameters used with NLStruct
Table 10: Hyperparameters used with NLStruct.
Parameter Value Parameter Value
Base encoder
BERT-base-uncased
SciBERT_scivocab_uncased
Learning rate 1e-3
Sequence length 256 Dropout 0
Train iteration 4,000 Warmup ratio 0
Optimizer AdamW Random seed 1 - 8 - 22 - 42 -100